If partice is in three dimensional box then write wavefunction
1
Expert's answer
2020-11-02T09:25:45-0500

The wavefunction is


"\psi_{n_x,n_y,n_z}(x,y,z)=\psi_{n_x}(x)\psi_{n_y}(y)\psi_{n_z}(z),\\\space\\
\psi_{n_x,n_y,n_z}(x,y,z)=\sqrt{\frac{8}{L_xL_yL_z}}\cdot\text{sin}\frac{n_x\pi x}{L_x}\cdot\text{sin}\frac{n_y\pi y}{L_y}\cdot\text{sin}\frac{n_z\pi z}{L_z}"